# Samba: Simple Hybrid State Space Models for Efficient Unlimited Context Language Modeling

### v1

@@ 4 Analysis
5 Conclusion
Acknowledgement
Appendix A Implementation Details
@@
For the GLA layer in the Sliding GLA architecture, we use the number of heads $d_{m}/384$, a key expansion ratio of 0.5, and a value expansion ratio of 1. For the RetNet layer we use a number of head that is half of the number of attention query heads, key expansion ratio of 1 and value expansion ratio of 2. The GLA and RetNet implementations are from the Flash Linear Attention repository [^1] https://github.com/sustcsonglin/flash-linear-attention [YZ24]. We use the FlashAttention-based implementation for Self-Extend extrapolation [^2] https://github.com/datamllab/LongLM/blob/master/self_extend_patch/Llama.py. The Mamba 432M model has a model width of 1024 and the Mamba 1.3B model has a model width of 2048. All models trained on SlimPajama have the same training configurations and the MLP intermediate size as Samba, unless otherwise specified. The training infrastructure on SlimPajama is based on a modified version of the TinyLlama codebase [^3] https://github.com/jzhang38/TinyLlama.
@@
In the generation configurations for the downstream tasks, we use greedy decoding for GSM8K, and Nucleus Sampling [HBD+19] with a temperature of $\tau=0.2$ and top-$p$ = 0.95 for HumanEval. For MBPP and SQuAD, we set $\tau=0.01$ and top-$p$ = 0.95.
Appendix B Additional Experiment Results
We perform instruction tuning for both Mistral 1.6B and Samba 1.7B on Passkey Retrieval using document length 4096, where we generated the data on the fly through randomly sampling a 5-digit integer passkey value and a location/depth between zero and the document length to insert the passkey. The model is then asked to generate the passkey given the full document. We train both models using batch size 2048, 250 warm-up steps with a peak learning rate of $1e^{-4}$, and 0.1 weight decay with AdamW [LH18] optimizer. In both cases, the loss converges quickly in 100-200 steps. During the evaluation, we measure the overall average accuracies of the passkey retrieval at the document length of [4k, 8k, 16k, 32k, 64k, 128k, 256k], for each length we evaluate at 11 different depths of the document (from 0, 0.1, 0.2, … to 1.0). In addition, for each location of the passkey (depth) in the document, we evaluate the model with five different passkeys to measure accuracy. As seen in Figure 7, the average passkey retrieval accuracy for Samba 1.7B almost reaches 100% in around 150 steps, while the accuracy for Mistral 1.6B remains low, demonstrating the extrapolation ability of the Samba architecture.
Appendix C Details of Entropy Measurement
Appendix D Limitations
Although Samba demonstrates promising memory retrieval performance through instruction tuning, its pre-trained base model has retrieval performance similar to that of the SWA-based model, as shown in Figure 7. This opens up future direction on further improving the Samba’s retrieval ability without compromising its efficiency and extrapolation ability. In addition, the hybridization strategy of Samba is not consistently better than other alternatives in all tasks. As shown in  Table 2, Mamba-SWA-MLP shows improved performance on tasks such as WinoGrande, SIQA, and GSM8K. This gives us the potential to invest in a more sophisticated approach to perform input-dependent dynamic combinations of SWA-based and SSM-based models.

### v2

@@ Appendix D Additional Analyses
Appendix E Details of Entropy Measurement
Appendix F Details of Downstream Long-context Evaluation
We use the GovReport (Huang et al., 2021) and the SQuALITY (Wang et al., 2022) datasets from the ZeroSCROLLS (Shaham et al., 2023) benchmark to evaluate models’ long-context summarization capability in the real world. After tokenizing with the Phi3-mini-4k tokenizer, the average document length for the GovReport dataset is 11,533 tokens, with a median of 10,332, a minimum of 1,493, and a maximum of 40,592 tokens. For the SQuALITY dataset, the average sequence length is 7,974 tokens, with a median of 8,145, a minimum of 5,457, and a maximum of 10,757 tokens. For evaluation, we use greedy decoding for both tasks. A maximum generation length of 450 tokens is applied for GovReport and 600 for SQuALITY.
Appendix G Implementation Details
@@
For the GLA layer in the Sliding GLA architecture, we use the number of heads $d_{m}/384$, a key expansion ratio of 0.5, and a value expansion ratio of 1. For the RetNet layer we use a number of head that is half of the number of attention query heads, key expansion ratio of 1 and value expansion ratio of 2. The GLA and RetNet implementations are from the Flash Linear Attention (Yang & Zhang, 2024) repository [^3] https://github.com/sustcsonglin/flash-linear-attention . We use the FlashAttention-based implementation for Self-Extend extrapolation [^4] https://github.com/datamllab/LongLM/blob/master/self_extend_patch/Llama.py. The Mamba 432M model has a model width of 1024 and the Mamba 1.3B model has a model width of 2048. All models trained on SlimPajama have the same training configurations and the MLP intermediate size as Samba, unless otherwise specified. The training infrastructure on SlimPajama is based on a modified version of the TinyLlama codebase [^5] https://github.com/jzhang38/TinyLlama.
@@
In the generation configurations for the downstream tasks, we use greedy decoding for GSM8K, and Nucleus Sampling (Holtzman et al., 2019) with a temperature of $\tau=0.2$ and top-$p$ = 0.95 for HumanEval. For MBPP and SQuAD, we set $\tau=0.01$ and top-$p$ = 0.95.
Appendix H Limitations & Broader Impact
Although Samba demonstrates promising memory retrieval performance through instruction tuning, its pre-trained base model has retrieval performance similar to that of the SWA-based model, as shown in Figure 8. This opens up future direction on further improving the Samba’s retrieval ability without compromising its efficiency and extrapolation ability. In addition, the hybridization strategy of Samba is not consistently better than other alternatives in all tasks. As shown in  Table 2, Mamba-SWA-MLP shows improved performance on tasks such as WinoGrande, SIQA, and GSM8K. This gives us the potential to invest in a more sophisticated approach to perform input-dependent dynamic combinations of SWA-based and SSM-based models (Ren et al., 2023). With the improved short-context performance and the long-term memorization ability of linear complexity LLMs such as Samba, cost-effective applications can be developed for personalized learning and automated tutoring. Samba can also be used for emotional accompaniment. The efficiency of the Samba architecture can save inference energy costs for models deployed on the edges, resulting in greener and more sustainable AI applications.